# Deep Graph Anomaly Detection: A Survey and New Perspectives

### v1

@@ Abstract
Graph Anomaly Detection, Graph Neural Networks, Deep Learning, Anomaly Detection, Graph Representation Learning
1 Introduction
@@
Related surveys. There have been several reviews on anomaly detection in recent years, e.g., [2, 4, 103, 89, 81, 113], but most of them are focused on non-deep-learning-based methods for GAD [2, 4, 103], or on general data rather than graph data [89, 7]. The studies [81, 113] are on deep GAD, but the reviews are limited to a relatively narrow point of view. For example, Ma et al. [81] focus on task-specific discussions, with limited reviews on the technical development, while Liu et al. [64] and Tang et al. [113] focus on establishing a performance benchmark for unsupervised and supervised GAD methods respectively. Another related work is Liu et al. [71], but it is restricted to supervised imbalanced graph learning. Although these surveys provide useful guidelines for the development of methods for GAD, it is difficult to understand the technical insights of existing methods and their limitations in addressing some unique challenges in GAD.
@@
  • The survey provides important insights into the problem complexities and the resulting challenges that are unique for the task of GAD (Sec. 2).
  • We introduce a novel taxonomy of current deep GAD methods, which offers in-depth understanding of the methods from three technical design perspectives, including GNN backbone design, proxy task design, and graph anomaly measures (Sec. 3).
  • We then introduce 13 fine-grained method categories under these three perspectives to provide more in-depth insights into the model designs (i.e., key intuition, assumption, learning objectives, advantages and disadvantages) and their capabilities in addressing the unique challenges in GAD (Secs. 4, 5, and 6).
  • We further discuss multiple important future research directions that involve largely unsolved open problems in GAD. Solutions in these directions would open up new opportunities for addressing the unique challenges in GAD (Sec. 7).
@@ 2 Problems and Challenges in GAD
2.1 Major Problem Complexities
@@ 2.2 Major Challenges
  • C5. Robust and interpretable GAD. GAD in real applications needs to be robust against various adverse conditions, such as abnormality camouflage (P8) [25, 37, 113] and unknown anomaly contamination [99, 98]. Addressing the abnormality camouflage or anomaly contamination may require models that can capture subtle differences between normal graph instances and camouflaged instances, and complex relationships within the graph as well. Besides, an explanation of why a graph instances is detected as anomaly can be crucial for the utility of the predictions in real applications [62, 106, 106], but it is a largely unexplored area. For example, in bank fraud detection, it is essential to provide a comprehensive explanation of the detected fraudulent activity for facilitating the subsequent investigation, but it is challenging to link the fraud to specific attributes of particular transactions (nodes) and their relations (edges) at a specific time period.
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, , [Deep GAD, text width=2cm, for tree=fill=hidden-grey!20 [GNN Backbone Design (§4), text width=3.3cm, for tree=fill=hidden-orange!20 [Discriminative GNNs, text width=2.8cm, for tree=fill=hidden-orange!20 [Aggregation Mechanism, text width=1.8cm, for tree=fill=hidden-orange!20 [ CARE-GNN [25]; GraphConsis [73]; PCGNN [66]; NGS [100]; GHRN [36]; H2-FDetector [108]; BLS [22]; FRAUDRE [142]; GAGA [124]; GDN [35]; GmapAD [82]; MTIGATE [12]; HedGe [147]; PMP [165]; RAND [6]; iGAD[143], text width=5.0cm, node options=align=left] ] [Feature Transformation, text width=1.8cm, for tree=fill=hidden-orange!20 [ GDN [35]; AMNet [10]; BWGNN [114] SEC-GFD [133]; RQGNN [24]; SplitGNN [125]; SmoothGNN [23] , text width=5.0cm, node options=align=left ] ] ] [Generative GNNs, text width=2.8cm, for tree=fill=hidden-orange!20 [Feature
@@
Interpolation, text width=1.8cm, for tree=fill=hidden-orange!20 [GraphSMOTE [150]; GraphENS [94]; DAGAD [60]; AuGAN [159], text width=5.0cm, node options=align=left ] ] [Noise
@@
Perturbation, text width=1.8cm, for tree=fill=hidden-orange!20 [ GGAD [99]; SDGG [8]; GODM [65]; DIFFAD [80]; ConsisGAD [14] , text width=5.0cm, node options=align=left ] ] ] ] [Proxy Task Design (§5), text width=3.3cm, for tree=fill=hidden-yellow!20 [Graph Reconstruction, text width=2.8cm, for tree=fill=hidden-yellow!20 [ DOMINANT [18]; AnomalyDAE [31]; GUIDE [141]; HO-GAT [45]; SpecAE [58]; ComGA [77]; ALARM [96]; REMAD [146]; MSAD [52]; GAD-NR [104]; Netwalk [140]; MUL-GAD [74]; AdoNE [5]; ResGCN [95]; Sub-CR [145]; VGOD [49]; AANE [26]; STRIPE [63]; HimNet [84], text width=7.3cm, node options=align=left ] ] [Graph Contrastive Learning, text width=2.8cm, for tree=fill=hidden-yellow!20 [CoLA [69] ; SL-GAD [155]; GCCAD [13]; ANEMONE [51]; GADMSL [27] PREM [85]; GRADATE [27]; CONAD [135] ; FMGAD [132]; SIGNET [67]; MAG [75]; Sub-CR [145]; ARISE [28]; HCM-A [47]; OCGTL [101]; NLGAD [29]; ACT [120]; FedCAD [55], text width=7.3cm, node options=align=left ] ] [Graph Representation Distillation, text width=2.8cm, for tree=fill=hidden-yellow!20 [GlocalKD [79]; GLADST [59]; FGAD [9], text width=7.3cm, node options=align=left ] ] [Adversarial Graph Learning, text width=2.8cm, for tree=fill=hidden-yellow!20 [AEGIS [17]; GAAN [15]; CFAD [131]; GADY [76]; GGA [83] , text width=7.3cm, node options=align=left ] ] [Score Prediction, text width=2.8cm, for tree=fill=hidden-yellow!20 [Meta-GAD [21]; SAD [116]; WEDGE [157] , text width=7.3cm, node options=align=left ] ] ] [Graph Anomaly Measures (§6), text width=3.3cm, for tree=fill=hidden-blue!20 [One-class Distance, text width=2.8cm, for tree=fill=hidden-blue!20 [OCGNN [122]; AAGNN [160]; DOHSC [148]; OCGTL [101]; DeepSphere [115]; Netwalk [140]; HRGCN [56], text width=7.3cm, node options=align=left ] ] [Community Adherence, text width=2.8cm, for tree=fill=hidden-blue!20 [ MHGL [158]; Netwalk [140], text width=7.3cm, node options=align=left ] ] [Local Affinity, text width=2.8cm, for tree=fill=hidden-blue!20 [TAM [98]; CLAD [53]; PREM[85]; ARC [68], text width=7.3cm, node options=align=left ] ] [Graph Isolation, text width=2.8cm, for tree=fill=hidden-blue!20 [ DIF [134]; GCAD [164], text width=7.3cm, node options=align=left ] ] ] ]
@@
3 Categorization of Deep GAD
3.1 Preliminaries
@@
3.2 Categorization of Deep GAD Methods
In order to facilitate a comprehensive understanding of the research progress in GAD, we introduce a new taxonomy that categorizes current GAD methods into three main groups,including GNN backbone design, GAD proxy task design, and graph anomaly measures, depending on the insights offered by each method. This enables us to review the GAD methods from three different technical perspectives. To elaborate the insights in each perspective, we further categorize the methods into fine-grained 13 groups. An overview of the taxonomy is shown in Figure 1.
More specifically, general GNNs can not be directly applied to GAD due to the aforementioned problem complexities, and thus, there is a group of studies that focus on designing suitable GNN backbones for GAD. The design of the GNN backbones can be divided into discriminative GNNs and generative GNNs according to the improvement of different modules in GNNs. The second main category of methods is on the GAD models constructed by optimizing a diverse set of well-crafted learning objective functions to form a proxy task that can guide the GAD models to capture diverse graph anomaly/normal patterns without the need for ground-truth labels. This category of methods can be further divided into five subcategories based on the modeling in the proxy tasks. Lastly, there is a group of methods that build GAD models based on anomaly measures that are designed specifically for graph data. These methods can be further grouped into four subcategories depending on the type of graph anomaly measures used. A summarization of representative algorithms for each type of GAD approaches is presented in Table II in Appendix A.
4 GNN Backbone Design
@@ 4.1.1 Aggregation Mechanism
where ${\bf{h}}_{i}^{(l)}$ is the feature representation of instance $i$ in the $l$-th layer, $\sigma$ is an activation function, and ${\bf{W}}^{(l)}$ is the training parameters in the $l$-th layer. The graph instance $o_{i}$ is often set as a node $v_{i}$, and $\mathcal{N}_{i}$ is typically the 1-hop neighborhood of the node $v_{i}$. However, due to the oversmoothing representation issue (C1 in Sec. 2.2) , directly applying such neighborhood aggregation mechanism can largely reduce the discriminability of of graph anomalies, especially for those whose abnormal behaviors are subtle (C3 in Sec. 2.2). Therefore, a variety of methods were proposed to enforce distinguishable representations for normal and abnormal graph instances throughout a number of feature aggregation iterations. These methods can be summarized via the following principled framework:
@@
$\bullet$ Hard Edge Selection. Popular aggregation mechanisms in GNN methods are built upon a homophily assumption that connected nodes come from the same class. Thus, the existence of non-homophily edges (i.e., edges that connect nodes of different classes, also referred to as heterophily edges below) in a GAD dataset can greatly hinders the discriminability of the learned feature representations. As shown in Fig. 2(a), one popular strategy is to instantiate $\Phi\left(\mathcal{N}_{i}\right)$ that prune the heterophily edges w.r.t. the normal class, referred to as hard edge selection. Below we review the methods in this line.
In order to enhance the homophily relations, CARE-GNN [25] devises a label-aware similarity measure to find informative neighboring nodes during the aggregation where $\Phi\left(\mathcal{N}_{i}\right)$ is instantiated by a node selector that chooses the neighbors with high similarity. Moreover, a reinforcement learning module is also used in [25] to find the optimal amounts of neighbors to be selected. MITIGATE [12] implements $\Phi\left(\mathcal{N}_{i}\right)$ via a masked aggregation mechanism that utilizes the distance-based clustering algorithm to choose a subset of high-representative nodes, in which the nodes that are closest to the cluster centers are chosen. GmpaAD [82] takes a similar clustering-based approach as MITIGATE, but it uses a differential evolutionary algorithm to find the optimal mapping strategy and generate the representative nodes given the selected candidates from a clustering method. On the other hand, H2-FDetector [108] categorizes the edges into homophily and heterophily connections in the graph, and further designs a new information aggregation strategy ensure that the homophily connections propagate similar information while the heterophily connections propagate different information.
@@
$\bullet$ Edge Synthesis. The edge selection function $\Phi$ focuses on local structure only. Edge synthesis function $\Psi$ can often be used to complement $\Phi$ in capturing more global patterns for GAD, as demonstrated in Fig. 2(b). For example, GHRN [36] prunes the inter-class edges by emphasizing and delineating the high-frequency components of the graph. Apart from edge reduction, the indirect link between nodes can also be beneficial for the normality representation learning [36]. To this end, GHRN introduces a global node selector $\Psi(\mathcal{V})$ that chooses nodes beyond the neighboring nodes of the target node to introduce this information into the feature aggregation, which can be seen as an edge synthesizer that connects distant nodes to a target node. PCGNN [66] specifies $\Psi(\mathcal{V})$ using a subgraph construction method consisting of class-label-balanced node and edge samplers to tackle potential issues arising from the skewed class distribution. To this end, it incorporates the label information into the sampling process to choose nodes and edges for its sub-graph construction. A distance function is also used to simplify the neighborhood in the sub-graphs. NSReg [121] leverages a novel normal structure regularization method where the normal-node-oriented relation is used to enforce strong normality into the representation learning to avoid overfitting to the labeled abnormal nodes. The relation representations are generated through a learnable transformation that fuses the representations of relevant nodes, which are subsequently used to optimize the normal-node-oriented relation prediction and the representation learner.
@@ 4.1.2 Feature Transformation
The aforementioned filters can be challenged by heterophily graphs since homophily graphs are assumed in these filters. To tackle this challenge, SEC-GFD [133] employs a hybrid band-pass filter to partition the graph spectrum into hybrid frequency bands, while SplitGNN [125] combines a band-pass graph filter with a tunable Beta Wavelet GNN to address the heterophily issue in node representation learning for GAD (C1 in Sec. 2.2).
@@
4.2 Generative GNNs
Generative GNN-based methods focus on synthesizing new graph instances to augment the existing graph data for enhancing model training on the new graph for GAD. This approach is motivated by the problem complexities like the scarcity of graph anomaly instances and their large variations (P5 and P7 in Sec. 2.2). The underlying key idea in these GAD methods is to synthesize outliers that can simulate graph anomalies in some specific properties to provide pseudo anomaly information for training the GAD models. In general, these methods can be summarized by the following formulation:

### v2

@@ Abstract
Graph Anomaly Detection, Graph Neural Networks, Anomaly Detection, Graph Representation Learning
I Introduction
@@
Related surveys. There have been several reviews on anomaly detection in recent years, e.g., [2, 4, 106, 91, 82, 116], but most of them are focused on non-deep-learning-based methods for GAD [2, 4, 106], or on general data rather than graph data [91, 7]. The studies [82, 116, 65] are on deep GAD, but the reviews are limited to a relatively narrow point of view. For example, Ma et al. [82] focus on task-specific discussions, with limited reviews on the technical development, while Liu et al. [65] and Tang et al. [116] focus on establishing a performance benchmark for unsupervised and supervised GAD methods respectively. Although these surveys provide useful guidelines for the development of methods for GAD, it is difficult to understand the technical insights of existing methods and their limitations in addressing some unique challenges in GAD.
@@
  • The survey provides important insights into the problem complexities and the resulting challenges that are unique for the task of GAD (Sec. II).
  • We introduce a novel taxonomy of current deep GAD methods, which offers in-depth understanding of the methods from three technical design perspectives, including GNN backbone design, proxy task design, and graph anomaly measures (Sec. III).
  • We then introduce 13 fine-grained method categories under these three perspectives to provide more in-depth insights into the model designs (i.e., key intuition, assumption, learning objectives, advantages and disadvantages) and their capabilities in addressing the unique challenges in GAD (Secs. IV, V, and VI).
  • We further discuss multiple important future research directions that involve largely unsolved open problems in GAD. Solutions in these directions would open up new opportunities for addressing the unique challenges in GAD (Sec. VII).
@@ II Problems and Challenges in GAD
II-A Major Problem Complexities
@@ II-B Major Challenges
  • C5. Robust and interpretable GAD. GAD in real applications needs to be robust against various adverse conditions, such as abnormality camouflage (P8) [24, 36, 116] and unknown anomaly contamination [102, 101]. Addressing the abnormality camouflage or anomaly contamination may require models that can capture subtle differences between normal graph instances and camouflaged instances, and complex relationships within the graph as well. Besides, an explanation of why a graph instance is detected as an anomaly can be crucial for the utility of the predictions in real applications [63, 109, 109], but it is a largely unexplored area. For example, in bank fraud detection, it is essential to provide a comprehensive explanation of the detected fraudulent activity for facilitating the subsequent investigation, but it is challenging to link the fraud to specific attributes of particular transactions (nodes) and their relations (edges) at a specific time period.
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, , [Deep GAD, text width=2cm, for tree=fill=hidden-grey!20 [GNN Backbone Design (§IV), text width=3.3cm, for tree=fill=hidden-orange!20 [Discriminative GNNs, text width=2.8cm, for tree=fill=hidden-orange!20 [Aggregation Mechanism, text width=1.8cm, for tree=fill=hidden-orange!20 [ CARE-GNN [24]; GraphConsis [74]; PCGNN [67]; NGS [103]; GHRN [35]; H2-FDetector [111]; BLS [21]; FRAUDRE [144]; GAGA [126]; GDN [34]; GmapAD [83]; MTIGATE [12]; HedGe [149]; PMP [167]; RAND [6]; iGAD[145]; GLHAD [38], text width=5.0cm, node options=align=left] ] [Feature Transformation, text width=1.8cm, for tree=fill=hidden-orange!20 [ GDN [34]; AMNet [10]; BWGNN [117] SEC-GFD [135]; RQGNN [23]; SplitGNN [127]; SmoothGNN [22] , text width=5.0cm, node options=align=left ] ] ] [Generative GNNs, text width=2.8cm, for tree=fill=hidden-orange!20 [Feature
@@
Interpolation, text width=1.8cm, for tree=fill=hidden-orange!20 [GraphSMOTE [152]; GraphENS [96]; DAGAD [61]; AuGAN [161]; gADAM [158], text width=5.0cm, node options=align=left ] ] [Noise
@@
Perturbation, text width=1.8cm, for tree=fill=hidden-orange!20 [ GGAD [102]; SDGG [8]; GODM [66]; DIFFAD [81]; ConsisGAD [14] , text width=5.0cm, node options=align=left ] ] ] ] [Proxy Task Design (§V), text width=3.3cm, for tree=fill=hidden-yellow!20 [Graph Reconstruction, text width=2.8cm, for tree=fill=hidden-yellow!20 [ DOMINANT [18]; AnomalyDAE [30]; GUIDE [143]; HO-GAT [45]; SpecAE [59]; ComGA [78]; ALARM [98]; REMAD [148]; MSAD [53]; GAD-NR [107]; Netwalk [142]; MUL-GAD [75]; AdoNE [5]; ResGCN [97]; Sub-CR [147]; VGOD [49]; AANE [25]; STRIPE [64]; HimNet [85]; SI-HGAD [168], text width=7.3cm, node options=align=left ] ] [Graph Contrastive Learning, text width=2.8cm, for tree=fill=hidden-yellow!20 [CoLA [70] ; SL-GAD [156]; GCCAD [13]; ANEMONE [52]; GADMSL [26] PREM [87]; GRADATE [26]; CONAD [137] ; FMGAD [134]; SIGNET [68]; MAG [76]; Sub-CR [147]; ARISE [27]; HCM-A [47]; OCGTL [104]; NLGAD [28]; ACT [122]; FedCAD [56], text width=7.3cm, node options=align=left ] ] [Graph Representation Distillation, text width=2.8cm, for tree=fill=hidden-yellow!20 [GlocalKD [80]; GLADST [60]; FGAD [9], text width=7.3cm, node options=align=left ] ] [Adversarial Graph Learning, text width=2.8cm, for tree=fill=hidden-yellow!20 [AEGIS [17]; GAAN [15]; CFAD [133]; GADY [77]; GGA [84] , text width=7.3cm, node options=align=left ] ] [Score Prediction, text width=2.8cm, for tree=fill=hidden-yellow!20 [Meta-GAD [20]; SAD [119]; WEDGE [159] , text width=7.3cm, node options=align=left ] ] ] [Graph Anomaly Measures (§VI), text width=3.3cm, for tree=fill=hidden-blue!20 [One-class Distance, text width=2.8cm, for tree=fill=hidden-blue!20 [OCGNN [124]; AAGNN [162]; DOHSC [150]; OCGTL [104]; DeepSphere [118]; Netwalk [142]; HRGCN [57], text width=7.3cm, node options=align=left ] ] [Community Adherence, text width=2.8cm, for tree=fill=hidden-blue!20 [ MHGL [160]; Netwalk [142], text width=7.3cm, node options=align=left ] ] [Local Affinity, text width=2.8cm, for tree=fill=hidden-blue!20 [TAM [101]; CLAD [54]; PREM[87]; ARC [69]; UNPrompt [86]; AnomalyGFM [100], text width=7.3cm, node options=align=left ] ] [Graph Isolation, text width=2.8cm, for tree=fill=hidden-blue!20 [ DIF [136]; GCAD [166], text width=7.3cm, node options=align=left ] ] ] ]
@@
III Categorization of Deep GAD
III-A Preliminaries
@@
III-B Categorization of Deep GAD Methods
In order to facilitate a comprehensive understanding of the research progress in GAD, we introduce a new taxonomy that categorizes current GAD methods into three main groups,including GNN backbone design, proxy GAD task design, and graph anomaly measures, depending on the insights offered by each method. This enables us to review the GAD methods from three different technical perspectives. To elaborate the insights in each perspective, we further categorize the methods into fine-grained 13 groups. An overview of the taxonomy is shown in Figure 1.
More specifically, general GNNs can not be directly applied to GAD due to the aforementioned problem complexities, and thus, there is a group of studies that focus on designing suitable GNN backbones for GAD. The design of the GNN backbones can be divided into discriminative GNNs and generative GNNs according to the improvement of different modules in GNNs. The second main category of methods is on the GAD models constructed by optimizing a diverse set of well-crafted learning objective functions to form a proxy task that can guide the GAD models to capture diverse graph anomaly/normal patterns without the need for ground-truth labels. This category of methods can be further divided into five subcategories based on the modeling in the proxy tasks. Lastly, there is a group of methods that build GAD models based on anomaly measures that are designed specifically for graph data. These methods can be further grouped into four subcategories depending on the type of graph anomaly measures used. A summarization of representative algorithms for each type of GAD approaches is presented in Table 1 in Appendix A.
IV GNN Backbone Design
@@ IV-A1 Aggregation Mechanism
where ${\bf{h}}_{i}^{(l)}$ is the feature representation of instance $i$ in the $l$-th layer, $\sigma$ is an activation function, and ${\bf{W}}^{(l)}$ is the training parameters in the $l$-th layer. The graph instance $o_{i}$ is often set as a node $v_{i}$, and $\mathcal{N}_{i}$ is typically the 1-hop neighborhood of the node $v_{i}$. $\text{AGG}(\cdot)$ is implemented using sum or mean aggregation, which sums or averages the representations of neighbor features. More advanced aggregation methods, such as attention-based aggregation [30] and LSTM [154], are also widely used. However, due to the oversmoothing representation issue (C1 in Sec. II-B) , directly applying such neighborhood aggregation mechanism can largely reduce the discriminability of of graph anomalies, especially for those whose abnormal behaviors are subtle (C3 in Sec. II-B). Therefore, a variety of methods were proposed to enforce distinguishable representations for normal and abnormal graph instances throughout a number of feature aggregation iterations. These methods can be summarized via the following principled framework:
@@
$\bullet$ Hard Edge Selection. Popular aggregation mechanisms in GNN methods are built upon a homophily assumption that connected nodes come from the same class.
Thus, the existence of non-homophily edges (i.e., edges that connect nodes of different classes, also referred to as heterophily edges below) in a GAD dataset can greatly hinders the discriminability of the learned feature representations. As shown in Fig. 2(a), one popular strategy is to instantiate $\Phi\left(\mathcal{N}_{i}\right)$ that prune the heterophily edges w.r.t. the normal class, referred to as hard edge selection. Below we review the methods in this line.
In order to enhance the homophily relations, CARE-GNN [24] devises a label-aware similarity measure to find informative neighboring nodes during the aggregation where $\Phi\left(\mathcal{N}_{i}\right)$ is instantiated by a node selector that chooses the neighbors with high similarity. Moreover, a reinforcement learning module is also used in [24] to find the optimal amounts of neighbors to be selected. MITIGATE [12] implements $\Phi\left(\mathcal{N}_{i}\right)$ via a masked aggregation mechanism that utilizes the distance-based clustering algorithm to choose a subset of high-representative nodes, in which the nodes that are closest to the cluster centers are chosen. GmpaAD [83] takes a similar clustering-based approach as MITIGATE, but it uses a differential evolutionary algorithm to find the optimal mapping strategy and generate the representative nodes given the selected candidates from a clustering method. On the other hand, H2-FDetector [111] categorizes the edges into homophily and heterophily connections in the graph, and further designs a new information aggregation strategy to ensure that the homophily connections propagate similar information while the heterophily connections propagate different information.
@@
$\bullet$ Edge Synthesis. The edge selection function $\Phi$ focuses on local structure only. Edge synthesis function $\Psi$ can often be used to complement $\Phi$ in capturing more global patterns for GAD, as demonstrated in Fig. 2(b). For example, GHRN [35] prunes the inter-class edges by emphasizing and delineating the high-frequency components of the graph. Apart from edge reduction, the indirect link between nodes can also be beneficial for the normality representation learning [35]. To this end, GHRN introduces a global node selector $\Psi(\mathcal{V})$ that chooses nodes beyond the neighboring nodes of the target node to introduce this information into the feature aggregation, which can be seen as an edge synthesizer that connects distant nodes to a target node. PCGNN [67] specifies $\Psi(\mathcal{V})$ using a subgraph construction method consisting of class-label-balanced node and edge samplers to tackle potential issues arising from the skewed class distribution. To this end, it incorporates the label information into the sampling process to choose nodes and edges for its sub-graph construction. A distance function is also used to simplify the neighborhood in the sub-graphs. NSReg [123] leverages a novel normal structure regularization method where the normal-node-oriented relation is used to enforce strong normality into the representation learning to avoid overfitting to the labeled abnormal nodes.
The relation representations are generated through a learnable transformation that fuses the representations of relevant nodes, which are subsequently used to optimize the normal-node-oriented relation prediction and the representation learner.
@@ IV-A2 Feature Transformation
The aforementioned filters can be challenged by heterophily graphs since homophily graphs are assumed in these filters. To tackle this challenge, SEC-GFD [135] employs a hybrid band-pass filter to partition the graph spectrum into hybrid frequency bands, while SplitGNN [127] combines a band-pass graph filter with a tunable Beta Wavelet GNN to address the heterophily issue in node representation learning for GAD (C1 in Sec. II-B).
@@
IV-B Generative GNNs
Generative GNN-based methods focus on synthesizing new graph instances to augment the existing graph data for enhancing model training on the new graph for GAD. This approach is motivated by the problem complexities like the scarcity of graph anomaly instances and their large variations (P5 and P7 in Sec. II-B). The underlying key idea in these GAD methods is to synthesize outliers that can simulate graph anomalies in some specific properties to provide pseudo anomaly information for training the GAD models. In general, these methods can be summarized by the following formulation: